answersLogoWhite

0

24.5

User Avatar

Wiki User

∙ 14y ago
User Avatar

What else can I help you with?

Related Questions

What is the probability of selecting two ribbons that are not blue if there are 7 yellow 6 blue 9 red and 3 green ribbons in a drawer?

The probability is 19/25 * 18/24 = 0.57


What is probability of throwing 9 heads in a row?

If you toss a coin often enough the probability is 1. The probability of 9 H in the first 9 tosses is (1/2)9 = 1/512


What is the probability of rolling a sum of 9 with two die?

The probability is 4/36 = 1/9


What is the probability of picking a 9 in a deck of cards?

The probability that a single card, drawn at random, is a 9 is 1/13.


What is the probability of landing on an odd number after spinning a spinner with 9 equal sectors numbered 1 through 9?

The probability is 5/9.


What is the probability 24 out of 32 as a decimal?

It is: 0.75


What is the probability of drawing a 9?

The answer depends on drawing a 9 from WHAT!


What is probability of drawing a seven or a nine out of cards?

Probability of drawing a 7 or 9 is the probability of draw a 7 plus the probability of draw a 9 which is 4/52 + 4/52 = 8/52 = 2/13 = 0.1538.


When two dice are thrown find probability of getting a sum of 8 or 10?

The probability is 8/36 = 4/9The probability is 8/36 = 4/9The probability is 8/36 = 4/9The probability is 8/36 = 4/9


What are odds of probability of 9 over 11?

For 9 to 11 odds of winning;Probability of winning:0.55, or;Chance of winning:55%


What is a square number less than 16 between 1-40 with a probability as a fraction?

9 is a square number less than 16. The number 9, in itself, has no probability since there is no "experiment" defined.9 is a square number less than 16. The number 9, in itself, has no probability since there is no "experiment" defined.9 is a square number less than 16. The number 9, in itself, has no probability since there is no "experiment" defined.9 is a square number less than 16. The number 9, in itself, has no probability since there is no "experiment" defined.


What is the probability of scoring a 9 on 2 dice?

1/9

Trending Questions
What is 155016 rounded to in thousands? If a p-value is 0.75 it's very likely that null hypothesis is false? What is 1.5 in fraction form? I usually have a period for 3 days this month it was 24 hours then i had a brown spotting 7 days later should you take a pregnancy test? What does 5th percentile crash test dummy means? What does a large MAD tell you about a set of data? What is statistical anomaly? What step in the composite risk management is focused on the probability and determining the severity of a hazard occurring? What expression using 1234 only once to equal 34? What is 532890 rounded to the nearest ten thousand? What is the only variable in the Poisson probability formula? What is the Difference between left hand gear box and right hand gear box? How many 4 digit even numbers are there if numbers can be repeated? What word means almost the same as guaranteed? When is it appropriate to use a normal model to model a set of data? Is it possible to get pregnant 2 weeks or 1 week before your period is suppossed to come And if you were would you miss your period because it was so close? What is the probability that in this family the first child born is a boy and the last child born is a girl of four children? A six-sided die numbered 1-6 is rolled and a card is drawn from a regular deck of 52 cards What is the probability of rolling a six and drawing a card with a 10 on it? What of the following numbers could not be a probability explain. 13 0 89 1 54? If you were testing how the number of batteries affect bulb brightness what heads would you use on a data table?